# Planting and MCMC Sampling from the Potts model

### v1

@@ 5. Proof for WSM within the ordered regime
the event that $v_{0},v_{1},\ldots,v_{\ell}$ form a path in $\hat{\boldsymbol{G}}$ (in this order) using the half-edges prescribed by $\xi$ and that they belong to the colour classes that $\tau$ prescribes. We will show later that [^2] We use the notation $f(n)\sim g(n)$ to denote that $\tfrac{f(n)}{g(n)}\rightarrow 1$ as $n\rightarrow\infty$.
Fix now an arbitrary vertex $v$ and let $P$ be an arbitrary sequence $(v_{0},v_{1},\ldots,v_{\ell})$ with $v_{0}=v$, together with a $q$-colouring $\tau=(\tau_{0},\ldots,\tau_{\ell})$ and pairing $\xi$ (as above). Condition on the event ${\mathcal{E}}_{P,\tau,\xi}$. Let $\boldsymbol{F}$ be the edge set obtained by keeping each monochromatic edge with probability $p$, and $\boldsymbol{F}_{v}$ be the set of these edges that are incident to $v$. Then, for the graph $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F}):=(\boldsymbol{V}\backslash\{v\},\boldsymbol{F}\backslash\boldsymbol{F}_{v})$, we will also show later that
$\mathbb{P}\Big{[}\begin{array}[]{c}\mbox{the largest component of $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$}\\ \mbox{avoids the $\tau$-coloured path $P$}\end{array}\,\Big{|}\,{\mathcal{E}}_{P,\tau,\xi}\Big{]}\leq(1+\tfrac{\varepsilon}{3})^{\ell}\hat{\varphi}^{\ell_{1}(\tau)}_{1},$ (5.4)
@@
$\mathbb{P}\Big{[}\begin{array}[]{c}\mbox{the largest component of $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$ avoids}\\ \mbox{some $\tau$-coloured path from~{}$v$ in~{}$G$}\end{array}\Big{]}\leq(1+\tfrac{\varepsilon}{3})^{\ell}d(d-1)^{\ell-1}\nu_{\mathrm{o},\tau_{0}}\prod^{\ell-1}_{k=0}\frac{\rho_{\mathrm{o},\tau_{k}\tau_{k+1}}}{\nu_{\mathrm{o},\tau_{k}}}\hat{\varphi}_{1}^{\mathbf{1}\{\tau_{k}=1\}}.$ (5.5)
At this stage, it remains to take a union bound over all possible colourings $\tau$ on $\ell+1$ vertices, i.e., sum the right-hand-side of (5.5) over all $\tau$. The resulting expression can be written more succinctly in the form $(1+\tfrac{\varepsilon}{3})d(d-1)^{l-1}\cdot\nu_{\mathrm{o}}^{T}M^{\ell}1$ where $M$ is the $q\times q$ matrix whose $(i,j)$-entry is given by $\frac{\rho_{\mathrm{o},ij}}{\hat{\nu}_{\mathrm{o},i}}$ and $\hat{\nu}_{\mathrm{o}}$ is the vector whose $i$-th entry is equal to $\nu_{\mathrm{o},1}/\hat{\varphi}_{1}$ if $i=1$ and $\nu_{\mathrm{o},i}$ if $i\in\{2,\ldots,q\}$. Note that $M=D^{-1/2}\hat{M}D^{1/2}$ where $D$ is the $q\times q$ diagonal matrix with $\hat{\nu}_{\mathrm{o}}$ on the diagonal and $\hat{M}$ is the symmetric matrix $\frac{\rho_{\mathrm{o},ij}}{(\hat{\nu}_{\mathrm{o},i}\hat{\nu}_{\mathrm{o},j})^{1/2}}$. In particular, we have that $\hat{M}$ and $M$ have the same eigenvalues, which are all real. Let $\sqrt{\hat{\nu}_{\mathrm{o}}}$ denote the vector whose entries are the square roots of the entries of $\hat{\nu}_{\mathrm{o}}$ Then, we can write
$\hat{\nu}_{\mathrm{o}}^{T}M^{\ell}1=\sqrt{\hat{\nu}_{\mathrm{o}}}^{T}\hat{M}^{\ell}\sqrt{\hat{\nu}_{\mathrm{o}}}\leq\tfrac{1}{\hat{\varphi}_{1}}\sqrt{\nu_{\mathrm{o}}}^{T}\hat{M}^{\ell}\sqrt{\nu_{\mathrm{o}}}$
@@
This finishes the proof of $\hat{M}\sqrt{\nu_{\mathrm{o}}}\leq A\sqrt{\nu_{\mathrm{o}}}$, and therefore we can bound the left-hand-side of (5.5) with $d(d-1)^{\ell-1}(A+\tfrac{\varepsilon}{2})^{\ell}$. Noting that the event in (5.5) is just $\mathbb{P}\Big{[}\boldsymbol{F}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big{)}\in\bigcup_{v\in V(G)}\mathcal{A}_{v}\Big{]}$, (5.2) follows by a union bound over $v$.
It remains to prove (5.3) and (5.4). We start with (5.3). Recall that the graph $\hat{\boldsymbol{G}}$ has $n_{i}=n\nu_{i}$ vertices of colour $i\in[q]$ and the number of edges between colour classes $i$ and $j$ with $i\neq j$ is given by $2m_{ij}:=dn\rho_{ij}$ if $i\neq j$, and $m_{ii}:=\frac{dn}{2}\rho_{ii}$ if $i=j$, cf. (2.4); moreover, the graph $\hat{\boldsymbol{G}}$ is uniformly distributed conditioned on these statistics. In particular, the number of ways to match the half-edges consistent with the vertex statistics $\{n_{i}\}$ and edge-statistics $\{m_{ij}\}$ is given by
We have
$\frac{\mathcal{N}^{\prime}}{\mathcal{N}}=\frac{(n-(\ell+1))!}{n!}\prod_{i\in[q]}\frac{n_{i}!}{(n_{i}-\ell_{i})!}\,\prod_{i\in[q]}\frac{(M_{i})!}{(dn_{i})!}\,\prod_{i\in[q]}\frac{2^{m_{ii}}(m_{ii})!}{2^{m_{ii}^{\prime}}(m_{ii}^{\prime})!}\,\prod_{i<j}\frac{(2m_{ij})!}{(2m_{ij}^{\prime})!}.$ (5.6)
It is a matter of a few manipulations now to massage the last expression into (5.3).
We next show (5.4). Condition on ${\mathcal{E}}_{P,\tau,\xi}$ and on the set of edges $E_{v}$ in $\hat{\boldsymbol{G}}$ incident to $v$. Let $\mathbf{d}$ be the degree sequence of the graph $\hat{\boldsymbol{G}}\backslash v=(V\backslash v,E\backslash(E_{P}\cup E_{v}))$. For a colour $i\in[q]$, let $\boldsymbol{V}_{\!i}$ denote the $i$-th colour class of $\hat{\boldsymbol{G}}\backslash v$ under $\hat{\boldsymbol{\sigma}}_{\mathrm{o}}$, $\mathbf{d}_{i}$ be the subsequence of $\mathbf{d}$ corresponding to vertices in $\boldsymbol{V}_{\!i}$, $\boldsymbol{E}_{i}$ be the set of monochromatic edges within $\boldsymbol{V}_{\!i}$, and $\boldsymbol{F}_{i}=\boldsymbol{F}\cap\boldsymbol{E}_{i}$. Conditioned on $|\boldsymbol{F}_{i}|=f_{i}$, the graph $\boldsymbol{G}_{i}=\big{(}\boldsymbol{V}_{\!i},\boldsymbol{F}_{i}\backslash(E_{P}\cup E_{v})\big{)}$ has the same distribution as the exact-edge model $\tilde{\boldsymbol{G}}_{n_{i},\mathbf{d}_{i},f_{i}}$ where $n_{i}=|\boldsymbol{V}_{\!i}|=n\nu_{\mathrm{o},i}+o(n)$. Note that for all $i\in[q]$, there are at most $\ell+1+d=O(n^{1/6})$ vertices in $\boldsymbol{d}_{i}$ with degree $\leq d$ (all other vertices have degree $d$).
Since $|\boldsymbol{E}_{i}|=\frac{dn}{2}\rho_{\mathrm{o},ii}+o(n)$, by a Chernoff bound we have that $|\boldsymbol{F}_{i}|=p\frac{dn}{2}\rho_{\mathrm{o},ii}+o(n)$ with probability $1-\mathrm{e}^{-\Omega(n)}$. It follows from Lemma 5.2, applied with densities $p_{i}=p\frac{\rho_{\mathrm{o},ii}}{\nu_{\mathrm{o},i}}$ for $i\in[q]$, that ${\boldsymbol{G}}_{1}$ is in the supercritical regime while ${\boldsymbol{G}}_{i}$ for $i\in\{2,\ldots,q\}$ is in the subcritical regime. Hence, it follows from Lemma 5.1 that w.p. $1-\mathrm{e}^{-\Omega(n)}$, the largest component ${\mathcal{C}}$ in $\boldsymbol{G}_{1}$ has size $\Omega(n)$ and the remaining components have all size $O(n^{2/3})$. Moreover, from Lemma 5.1 part (2) we obtain that [^4] Note that, because the edges in $E_{v}$ are excluded from $\hat{\boldsymbol{G}}\setminus v$, all but at most $d+2$ vertices of the path $P$ have degree $d-2$ in $\hat{\boldsymbol{G}}\setminus v$; the contribution of vertices with degree $<d-2$ is at most a constant which is absorbed into the $(1+\frac{\varepsilon}{3})^{\ell}$ factor in (5.7).
$\mathbb{P}[\,{\mathcal{C}}\mbox{ avoids the path }P\mid{\mathcal{E}}_{P,\tau,\xi},E_{v}]\leq(1+\tfrac{\varepsilon}{3})^{\ell}(p\varphi+1-\varphi)^{(d-2)\ell_{1}}\leq(1+\tfrac{\varepsilon}{3})^{\ell+1}\hat{\varphi}_{1}^{\ell_{1}}.$ (5.7)
By contrast, for $i\in\{2,\ldots,q-1\}$, w.p. $1-\mathrm{e}^{-\Omega(n)}$, the largest component in $\boldsymbol{G}_{i}$ has size $O(n^{1/2})$. Since the union of the $\boldsymbol{G}_{i}$’s over $i\in[q]$ is the graph $\widetilde{\boldsymbol{G}}_{P,v}=\big{(}\boldsymbol{V}\backslash\{v\},\boldsymbol{F}\backslash(E_{P}\cup E_{v})\big{)}$, by a union bound it follows that, w.p. $1-\mathrm{e}^{-\Omega(n)}$, ${\mathcal{C}}$ is the largest component in $\widetilde{\boldsymbol{G}}_{P,v}$ with $\Omega(n)$ size, while all the others have size $O(n^{2/3})$. Now add the monochromatic edges of the path $P$, each percolated with probability $p$, to obtain the graph $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})=(\boldsymbol{V}\backslash\{v\},\boldsymbol{F}\backslash\boldsymbol{F}_{v})$. Let ${\mathcal{C}}^{\prime}$ be the component of $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$ with ${\mathcal{C}}\subseteq{\mathcal{C}}^{\prime}$. Since we added only edges on the path $P$, we have
$\mathbb{P}[\,{\mathcal{C}}^{\prime}\mbox{ avoids the path }P\mid{\mathcal{E}}_{P,\tau,\xi}]=\mathbb{P}[\,{\mathcal{C}}\mbox{ avoids the path }P\mid{\mathcal{E}}_{P,\tau,\xi}].$ (5.8)
Moreover, the length of the path $P$ is $\leq n^{1/6}$, so $|{\mathcal{C}}^{\prime}|-|{\mathcal{C}}|=O(n^{5/6})$. Similarly, any component in $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$ other than ${\mathcal{C}}^{\prime}$ has size $O(n^{2/3})$, and in particular ${\mathcal{C}}^{\prime}$ is the largest component in $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$. Combining this with (5.7) and (5.8), we obtain (5.4). ∎
@@
It only remains now to bound the quantity $A$ appearing in Lemma 5.3 by a suitably small constant so that we can take a union bound over paths; it is for this step that we require $q\geq(5d)^{5}$. We show in particular the following.
@@
Let $d\geq 3$, $q\geq(5d)^{5}$ and $\beta\geq\beta_{c}$. Then, for all $i\in[q]$, it holds that ${\small(1-\sqrt{\hat{\varphi}_{1}})}\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}>1-\frac{1}{(d-1)^{5}}$.
Proof.
$\frac{\rho_{\mathrm{o},{i1}}}{\nu_{\mathrm{o},i}}=1-R\mbox{ where }R:=\tfrac{t^{d+1}+(q-2)t^{d}-(q-1)t}{(t^{d}-t)(t^{d-1}+q-1)}$ (5.9)
To upper bound $\hat{\varphi}_{1}$, we need first to find the solution of $\varphi_{1}=(p_{1}\varphi_{1}+1-p_{1})^{d-1}$ for $p_{1}=p\frac{\rho_{\mathrm{o},11}}{\nu_{\mathrm{o},1}}$. Similarly to above, we have $\rho_{\mathrm{o},11}=\frac{\mathrm{e}^{\beta}t^{2(d-1)}}{(t^{d-1}+q-1)^{2}+(\mathrm{e}^{\beta}-1)(t^{2d-2}+(q-1))}$, so using the expression for $\mathrm{e}^{\beta}-1$ from (2.1) we obtain after some algebra that $p_{1}=\frac{t^{d-1}(t-1)}{t^{d}-t}$ and $\varphi_{1}=1/t^{d-1}$. Since $t\geq t_{c}$ for $\beta\geq\beta_{c}$, we obtain that $\varphi_{1}\leq\frac{1}{(q-1)^{2(d-1)/d}}$. It follows that $\hat{\varphi}_{1}=(p_{1}\varphi_{1}+1-p_{1})^{d-2}=\varphi_{1}^{(d-2)/(d-1)}\leq\frac{1}{(q-1)^{2(d-2)/d}}$, so $\sqrt{\hat{\varphi}_{1}}\leq\frac{1}{(q-1)^{(d-2)/d}}$. For $q\geq(5d)^{5}$, numerical calculations give that this is less than $\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ (the inequality is tighter for small values of $d$). A similar argument shows that $R<\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ when $q\geq(5d)^{5}$; to see this, we have from the expression in (5.9) that $R\leq R^{\prime}$ where $R^{\prime}:=\tfrac{t+q-1}{t^{d-1}+q-1}$. We have that $R^{\prime}$ is decreasing in $t$, so using that $t\geq t_{c}$, we obtain that $R^{\prime}\leq\tfrac{t_{c}+q-1}{t^{d-1}_{c}+q-1}=\frac{1}{(q-1)^{(d-2)/d}}$; from the same numerical inequality as above, we therefore conclude that $R$ is less than $\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ when $q\geq(5d)^{5}$. Hence, $\frac{1}{(d-1)^{5}}>\sqrt{\hat{\varphi}_{1}}+R$, as needed. ∎
@@ 5.1. WSM within the ordered phase
Let $G=(V,E)$ be a graph, $v\in V$, $\ell\geq 1$ and let $F\subseteq E$ be a RC configuration. Say that $\mathcal{S}\subseteq V(B_{\ell}(v))\setminus\{v\}$ is a wired boundary around $v$ in $B_{\ell}(v)$, if
  • $\mathcal{S}$ is a cut set separating $v$ from the outside of the ball. That is, ${\mathcal{C}}_{v}$, the component of $v$ in $G\setminus\mathcal{S}$, is disjoint from $G\setminus B_{\ell}(v)$; and
  • in the graph $(V\setminus V({\mathcal{C}}_{v}),\{e\in F:e\cap V({\mathcal{C}}_{v})=\emptyset\})$, all vertices of $\mathcal{S}$ are in the same component.
@@
$\big{|}\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big{|}\leq\pi_{G,\mathrm{o}}(\forall\mathcal{S}.\text{ }\mathcal{S}\text{ is not a wired boundary around }v\text{ in }B_{\ell}(v))+\mathrm{e}^{-\Omega(n)}$
@@
We bound the total variation distance by a probability that two coupled configurations, $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$ and $\boldsymbol{F}^{+}\sim\pi_{G,v,\ell}^{+}$, disagree on edge $e$.
Consider first the outside of the ball $B_{\ell}(v)$. If at least $(n\mathfrak{m}_{\mathrm{o}}(\beta_{c})-n\rho/4$ (where $\rho=\mathfrak{m}_{\mathrm{o}}(\beta_{c})-\mathfrak{m}_{\mathrm{d}}(\beta_{c})$ as in Definition 3.1) edges outside the ball belong to some configuration, then regardless of the state of the edges in the ball, this configuration is ordered.
Let $\mathcal{O}$ be the event that $\boldsymbol{F}_{\mathrm{o}}$ contains at least $(n\mathfrak{m}_{\mathrm{o}}(\beta_{c})-n\rho/4$ edges outside of the ball, and note that conditional on $\mathcal{O}$ we can thus drop the conditioning on the ordered phase.
Note that $|E(B_{\ell}(v))|=O(n^{1/2})$, thus for $n$ large enough, $|E(B_{\ell}(v))|\leq n\rho/8$. Hence, for $n$ large enough, $\mathbb{P}(\overline{\mathcal{O}})\leq\mathbb{P}(|\boldsymbol{F}_{\mathrm{o}}|\leq n\mathfrak{m}_{\mathrm{o}}-n\rho/8)$, which is, by Corollary B.2, whp over $G$, bounded by $\mathrm{e}^{-\Omega(n)}$.
Now we can use monotonicity property of the random cluster model (see e.g. [RCrandom, Appendix B]), to get that $\pi_{G,v,\ell}^{+}$ stochastically dominates $\pi_{G,\mathrm{o}}(\cdot\mid\mathcal{O})$ on $B_{\ell}(v)$, hence conditional on $\mathcal{O}$ we can perfectly couple $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$ and $\boldsymbol{F}^{+}\sim\pi_{G,v,\ell}^{+}$ such that $\boldsymbol{F}_{\mathrm{o}}\cap E(B_{\ell}(v))\subseteq\boldsymbol{F}^{+}\cap E(B_{\ell}(v))$. Then we can bound the probability the two configurations disagree on $e$ by at most
Conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}},\mathcal{O}$, we can see $\boldsymbol{F}_{\mathrm{o}}$, $\boldsymbol{F}^{+}$ generated as first generating the edges outside of the component of $v$ in $G\setminus\mathcal{S}$ – call these edges $E_{O}$, and then, conditional on $E_{O}$ and ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$ the remainder of the edges. Note that for $\mathcal{S}\in\mathcal{S}_{B}$, the events “$\mathcal{S}$ is a wired boundary in $\boldsymbol{F}_{\mathrm{o}}$” (which implies it is a wired boundary also in $\boldsymbol{F}^{+}$) and $\mathcal{O}$ are determined by $\boldsymbol{F}_{\mathrm{o}}\cap E_{O}$.
Thus, conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$, $\mathcal{O}$, and then any (possible) subconfiguration of $(\boldsymbol{F}_{\mathrm{o}},\boldsymbol{F}^{+})$ on $E_{O}$, they both have a wired boundary $\mathcal{S}$. Then by [RCrandom, Observation 51], the marginals edges in $E\setminus E_{O}$ agree, hence since $e\not\in E_{O}$ and by the fact that $(\boldsymbol{F}_{\mathrm{o}},\boldsymbol{F}_{+})$ were perfectly coupled, the probability of disagreeing on $e$ conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$ and $\mathcal{O}$ is zero. ∎
Next we show that we can relate lack of long paths avoiding the largest component with the existence of the wired boundary.
Lemma 5.7.
Let $G=(V,E)$ be a graph, let $v\in V$ be an arbitrary vertex, and let $\ell\geq 1$ an integer. For any $F\subseteq E$, let $F_{v}$ be the set of edges in $F$ that are incident to $v$ and let $G_{v,F}:=(V\setminus\{v\},F\setminus F_{v})$. Suppose that all simple $-\ell$ paths from $v$ in $G$ intersect the largest component of $G_{v,F}$. Then $F$ has a wired boundary around $v$ in $B_{\ell}(v)$.
Proof.
Consider the set of paths $\mathcal{P}$ of length $\ell$ in $G$ that start in $v$.
Claim 2. No vertices of ${\mathcal{C}}_{1}$ belong to ${\mathcal{C}}_{v}$.
See 3.7
Proof.
By Lemma 5.4, for this parameter range we have $A:=\{\max_{i}(1-(1-\sqrt{\hat{\varphi}_{1}}))\frac{\rho_{o,i1}}{\nu_{o,i}}\}<(d-1)^{-5}$. Let $\varepsilon>0$ be a small constant (depending on $d$) such that $A+\varepsilon\leq(d-1)^{-(5+\varepsilon)}$. Let $\delta:=\frac{\varepsilon}{4(4+\varepsilon)}$. By Lemma 5.6, $\big{|}\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big{|}$ is at most
@@
$\mathrm{e}^{-\Omega(n)}+1-\pi_{G,\mathrm{o}}(\text{ all simple length-$\ell$ paths from $v$ in $G$ intersect the largest componentof $G_{v,F}$ }).$

### v2

@@ 5. Proof for WSM within the ordered regime
the event that $v_{0},v_{1},\ldots,v_{\ell}$ form a path in $\hat{\boldsymbol{G}}$ (in this order) using the half-edges prescribed by $\xi$ and that they belong to the colour classes that $\tau$ prescribes. In Appendix E.1, we give an exact expression for $\mathbb{P}({\mathcal{E}}_{P,\tau,\xi})$ in terms of $\nu$ and $\rho$ which yields that
$\mathbb{P}({\mathcal{E}}_{P,\tau,\xi})\sim(dn)^{-\ell}\nu_{\mathrm{o},\tau_{0}}\prod^{\ell-1}_{k=0}\frac{\rho_{\mathrm{o},\tau_{k}\tau_{k+1}}}{\nu_{\mathrm{o},\tau_{k}}},$ (5.3)
where $f(n)\sim g(n)$ denotes $\frac{f(n)}{g(n)}\rightarrow 1$ as $n\rightarrow\infty$.
[…]
@@
[…]
At this stage, it remains to take a union bound over all possible colourings $\tau$ on $\ell+1$ vertices, i.e., sum the right-hand-side of (5.5) over all $\tau$. The resulting expression can be written more succinctly in the form $(1+\tfrac{\varepsilon}{3})d(d-1)^{l-1}\cdot\nu_{\mathrm{o}}^{T}M^{\ell}1$ where $M$ is the $q\times q$ matrix whose $(i,j)$-entry is given by $\frac{\rho_{\mathrm{o},ij}}{\hat{\nu}_{\mathrm{o},i}}$ and $\hat{\nu}_{\mathrm{o}}$ is the vector whose $i$-th entry is equal to $\nu_{\mathrm{o},1}/\hat{\varphi}_{1}$ if $i=1$ and $\nu_{\mathrm{o},i}$ if $i\in\{2,\ldots,q\}$. Note that $M=D^{-1/2}\hat{M}D^{1/2}$ where $D$ is the $q\times q$ diagonal matrix with $\hat{\nu}_{\mathrm{o}}$ on the diagonal and $\hat{M}$ is the symmetric matrix $\frac{\rho_{\mathrm{o},ij}}{(\hat{\nu}_{\mathrm{o},i}\hat{\nu}_{\mathrm{o},j})^{1/2}}$. In particular, we have that $\hat{M}$ and $M$ have the same eigenvalues, which are all real. Let $\sqrt{\hat{\nu}_{\mathrm{o}}}$ denote the vector whose entries are the square roots of the entries of $\hat{\nu}_{\mathrm{o}}$. Then, we can write
$\mbox{$\hat{\nu}_{\mathrm{o}}^{T}M^{\ell}1=\sqrt{\hat{\nu}_{\mathrm{o}}}^{T}\hat{M}^{\ell}\sqrt{\hat{\nu}_{\mathrm{o}}}\leq\tfrac{1}{\hat{\varphi}_{1}}\sqrt{\nu_{\mathrm{o}}}^{T}\hat{M}^{\ell}\sqrt{\nu_{\mathrm{o}}}$}.$
@@
This finishes the proof of $\hat{M}\sqrt{\nu_{\mathrm{o}}}\leq A\sqrt{\nu_{\mathrm{o}}}$, and therefore we can bound the left-hand-side of (5.5) with $d(d-1)^{\ell-1}(A+\tfrac{\varepsilon}{2})^{\ell}$. Noting that the event in (5.5) is just $\mathbb{P}\Big[\boldsymbol{F}\big(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big)\in\bigcup_{v\in V(G)}\mathcal{A}_{v}\Big]$, (5.2) follows by a union bound over $v$. ∎
@@
We next bound the quantity $A$ appearing in Lemma 5.3 by a suitably small constant so that we can take a union bound over paths; it is for this step that we require $q\geq(5d)^{5}$. The proof is in Appendix E.2.
@@
Let $d\geq 3$, $q\geq(5d)^{5}$ and $\beta\geq\beta_{c}$. Then, for all $i\in[q]$, it holds that ${(1-\sqrt{\hat{\varphi}_{1}})}\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}>1-\frac{1}{(d-1)^{5}}$.
@@ 5.1. WSM within the ordered phase
Let $G=(V,E)$ be a graph, $v\in V$, $\ell\geq 1$ and let $F\subseteq E$ be a RC configuration. Consider the $\ell$-neighbourhood $B_{\ell}(v)$ of $v$ in $G$. We say that $S\subseteq V(B_{\ell}(v))\setminus\{v\}$ is a wired boundary around $v$ in $B_{\ell}(v)$, if
  • $S$ is a cut set separating $v$ from the outside of the ball. That is, ${\mathcal{C}}_{v}$, the component of $v$ in $G\setminus S$, is disjoint from $G\setminus B_{\ell}(v)$; and
  • in the graph $(V\setminus V({\mathcal{C}}_{v}),\{e\in F:e\cap V({\mathcal{C}}_{v})=\emptyset\})$, all vertices of $S$ are in the same component.
@@
$\big|\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big|\leq\pi_{G,\mathrm{o}}\big(\mbox{\emph{no wired boundary around $v$ in $B_{\ell}(v)$}}\big)+\mathrm{e}^{-\Omega(n)}.$
To prove the lemma, we compare with $\pi_{G}$. When $\beta>\beta_{c}$, we obtain, from Corollary 3.2, that $\left\|\pi_{G,\mathrm{o}}-\pi_{G}\right\|_{\mathrm{TV}}=\mathrm{e}^{-\Omega(n)}$. Then we can bound $|\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G}(1_{e})|$ by the probability (under $\pi_{G}$) that there is no wired boundary around $v$ in $B_{\ell}(v)$: informally, this is because probability of $1_{e}$ conditional on the wired boundary is the same under both $\pi_{G}$ and $\pi_{G,v,\ell}^{+}$, and due to the monotonicity of the RC distribution, the probability of the wired boundary under $\pi_{G,v,\ell}^{+}$ is at least that under $\pi_{G}$. For more details as well as the case when $\beta=\beta_{c}$ see the full proof of the Lemma is in the Appendix E.3.
By Lemma 5.3 all simple length-$l$ paths from $v$ intersect the largest component of $G_{v,F}$, so we can construct a wired boundary $S$ by taking, for each length-$\ell$ path from $v$ in $G$, the first vertex on the path belonging to the largest component of $G_{v,F}$. We thus show the following lemma in Appendix E.3.
Lemma 5.7.
Let $G=(V,E)$ be a graph, let $v\in V$ be an arbitrary vertex, and let $\ell\geq 1$ an integer. For any $F\subseteq E$, let $F_{v}$ be the set of edges in $F$ that are incident to $v$ and let $G_{v,F}:=(V\setminus\{v\},F\setminus F_{v})$. Suppose that all simple length-$\ell$ paths from $v$ in $G$ intersect the largest component of $G_{v,F}$. Then $F$ has a wired boundary around $v$ in $B_{\ell}(v)$.
From these two lemmas we can now obtain WSM within the ordered phase. See 3.7
@@
By Lemma 5.4, we have $A:=\{\max_{i}(1-(1-\sqrt{\hat{\varphi}_{1}}))\frac{\rho_{o,i1}}{\nu_{o,i}}\}<(d-1)^{-5}$. Let $\varepsilon>0$ be a small constant (depending on $d$) such that $A+\varepsilon\leq(d-1)^{-(5+\varepsilon)}$. Let $\delta:=\frac{\varepsilon}{4(4+\varepsilon)}$. By Lemma 5.6, $\big|\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big|$ is at most
@@
[…]